OpenStudy (anonymous):

Let X be the number of successes in six independent trials of a binomial experiment in which the probability of success is p = 2/5 Find the following probabilities. (Round your answers to four decimal places.) (a) P(X = 5) (b) P(3 ≤ X ≤ 5)

OpenStudy (anonymous):

This is plug and chug.

OpenStudy (anonymous):

\[ \mathcal P(X=k) = {n\choose k}p^k(1-p)^{n-k}={6\choose k}\left(\frac 25\right)^k\left(\frac 35\right)^{6-k} \]

OpenStudy (anonymous):

\[ \mathcal P(a\leq X\leq b) = \sum_{k=a}^{b}P(X=k) = \sum_{k=a}^{b}{6\choose k}\left(\frac 25\right)^k\left(\frac 35\right)^{6-k} \]